Let D={(x,y,z)∈R^3:−1≤x≤2,0≤y≤1,0≤z≤π}. The answer to ∭D yzdV

is
Select one:
a. (3π)/4

b. 3π^2

c. 3/4

d. 75π^2


e. (3π^2)/4
1
Expert's answer
2020-06-04T19:29:51-0400

"\begin{aligned}
\iiint_{D} Y Z d v &=\int_{0}^{\pi} \int_{0}^{1} \int_{-1}^{2} y z d x d y d z \\
&=\int_{0}^{\pi} \int_{-1}^{2}\left(\int_{0}^{1} y z d y\right) d x d z \\
&=\int_{0}^{\pi} \int_{-1}^{2}\left(\left.\frac{1}{2} y^{2} z\right|_{0} ^{1}\right) d x d z \\
&=\int_{0}^{\pi} \int_{-1}^{2}\left(\frac{1}{2} z\left(1^{2}-0^{2}\right)\right) d x d z \\
&=\int_{-1}^{2}\left(\int_{0}^{\pi} \frac{1}{2} z d z\right) d x \\
&=\int_{-1}^{2}\left(\frac{1}{2}\left(\left.\frac{1}{2} z^{2}\right|_{0} ^{2}\right) d x\right.\\
&=\int_{-1}^{2} \frac{1}{4} \pi^{2} d x \\
&=\left.\frac{1}{4} \pi^{2} x\right|_{-1} ^{2} \\
&=\frac{1}{4} \pi^{2}(2-(-1)) \\
&=\frac{3}{4} \pi^{2} \\
&=\operatorname{The} \operatorname{answer} \operatorname{is} \left(\frac{3 \pi^{2}}{4}\right)
\end{aligned}"